OpenStudy (mtalhahassan2):

Need help!! evaluate the following logarithms a) log base 25 5^3

Nnesha (nnesha):

power rule \[\huge\rm log_b x^y = y \log_b x\] then use change of base formula \[\huge\rm log_a b = \frac{ \log b }{ \log a }\]

OpenStudy (mtalhahassan2):



OpenStudy (mtalhahassan2):

like that

Nnesha (nnesha):

what is the exponent of 5 ?

OpenStudy (mtalhahassan2):

1

Nnesha (nnesha):

:o exponent of 5 in the original question

OpenStudy (mtalhahassan2):